# Learning Heuristics for Transit Network Design and Improvement with Deep Reinforcement Learning

### v1

@@ 3.1 Markov Decision Process Formulation
A Markov Decision Process (MDP) is a formalism commonly used to define problems in RL (Sutton and Barto, 2018, Chapter 3). In an MDP, an agent interacts with an environment over a series of timesteps. At each timestep $t$, the environment is in a state $s_{t}\in\mathcal{S}$, and the agent takes some action $a_{t}\in\mathcal{A}_{t}$, where $\mathcal{A}_{t}$ is the set of available actions at $t$. The environment then transitions to a new state $s_{t+1}\in\mathcal{S}$ according to the state transition distribution $P(s_{t+1}|s_{t},a_{t})$, and the agent receives a numerical reward $R_{t}\in\mathbb{R}$ according to the reward distribution $P(R_{t}|s_{t},a_{t},s_{t+1})$. The agent chooses actions according to its policy $\pi(a_{t}|s_{t})$, which is a probability distribution over $\mathcal{A}_{t}$ given $s_{t}$ in each state. In RL, the goal is to learn a policy $\pi$ that maximizes the return $G_{t}$, defined as a time-discounted sum of rewards:
$G_{t}=\sum^{t_{end}}_{t^{\prime}=t}\gamma^{t^{\prime}-t}R_{t^{\prime}}$ (2)
Where $\gamma\in[0,1]$ is a parameter that discounts rewards farther in the future, and $t_{end}$ is the final time-step of the MDP. The sequence of states visited, actions taken, and rewards received from $t=0$ to $t=t_{end}$ constitutes one episode.
We here describe the MDP we use to represent a construction approach to the Transit Network Design Problem. As shown in Equation 3, the state $s_{t}$ is composed of the set of routes $\mathcal{R}_{t}$ planned so far, and an incomplete route $r_{t}$ which is being planned.
@@
The starting state is $s_{0}=(\mathcal{R}_{0}=\{\},r_{0}=[])$. At a high level, the MDP alternates at every timestep between two modes: on odd-numbered timesteps, the agent selects an extension to the route $r_{t}$ that it is currently planning; on even-numbered timesteps, the agent chooses whether to stop extending $r_{t}$ and add it to $\mathcal{R}_{t}$.
On odd-numbered timesteps, the available actions are drawn from SP, the set of shortest paths between all node pairs. If $r_{t}=[]$, then $\mathcal{A}_{t}=\{a\;|\;a\in\textup{SP},|a|\leq MAX\}$. Otherwise, $\mathcal{A}_{t}$ is comprised of paths $a\in\textup{SP}$ satisfying all of the following conditions:
@@
Once a path $a_{t}\in\mathcal{A}_{t}$ is chosen, $r_{t+1}$ is formed by appending $a_{t}$ to the beginning or end of $r_{t}$ as appropriate: $r_{t+1}=\textup{combine}(r_{t},a_{t})$.
@@
$\mathcal{A}_{t}=\begin{cases}\{\textup{continue}\}&\text{if}|r_{t}|<MIN\\ \{\textup{halt}\}&\text{if}|r_{t}|=MAX\\ \{\textup{continue},\textup{halt}\}&\text{otherwise}\end{cases}$ (4)
If $a_{t}=\textup{halt}$, $r_{t}$ is added to $\mathcal{R}_{t}$ to get $\mathcal{R}_{t+1}$, and $r_{t+1}=[]$ is a new empty route. If $a_{t}=\textup{continue}$, then $\mathcal{R}_{t+1}$ and $r_{t+1}$ are unchanged from timestep $t$.
When $|\mathcal{R}_{t}|=S$, the episode terminates, giving the final reward $R_{t}=-C(\mathcal{C},\mathcal{R}_{t})$. At all prior steps, $R_{t}=0$.
This MDP formalization imposes some helpful biases on the solution space. First, it requires any route connecting $i$ and $j$ to stop at all nodes along some path between $i$ and $j$, biasing planned routes towards covering more nodes. Second, it biases routes towards directness by forcing them to be composed of shortest paths. While policy may construct arbitrarily indirect routes by choosing paths with length 2 at every step, this is unlikely because in a realistic street graph, the majority of paths in SP are longer than one edge. Third, the alternation between deciding to continue or halt a route and deciding how to extend the route means that the probability of halting does not depend on how many different extensions are available.
@@ 3.2 Cost Function
Where $\tau_{\mathcal{R}ij}$ is the time of the shortest transit trip from $i$ to $j$ given $\mathcal{R}$, including a time penalty $p_{T}$ for each transfer.
@@
To enforce the constraints on $\mathcal{R}$, we use a third term $C_{c}$, which is the fraction of node pairs that are not connected by $\mathcal{R}$ plus a measure of how much $|r|>MAX$ or $|r|<MIN$ across all routes. The cost function is then:
@@
The weight $\alpha\in[0,1]$ controls the trade-off between passenger and operating costs, while $\beta$ is the penalty assigned for each cost violation. $w_{p}$ and $w_{o}$ are re-scaling constants chosen so that $w_{p}C_{p}$ and $w_{o}C_{o}$ both vary roughly over the range $[0,1)$ for different $\mathcal{C}$ and $\mathcal{R}$; this is done so that $\alpha$ will properly balance the two, and to stabilize training of the GNN policy. The values used are $w_{p}=(\max_{i,j}T_{ij})^{-1}$ and $w_{o}=(S\max_{i,j}T_{ij})^{-1}$, where $T$ is an $n\times n$ matrix of shortest-path driving times between every node pair.
@@ 4.1 Learned Constructor
In order to learn heuristics for transit network design, we first train a neural net policy $\pi_{\theta}(a|s)$, parameterized by $\theta$, to maximize the cumulative return $G_{t}$ on the construction MDP described in subsection 3.1. By then following this policy on the MDP for some city $\mathcal{C}$, randomly sampling $a_{t}$ at each step with probabilities given by $\pi(\cdot|s_{t})$, we can obtain a transit network $\mathcal{R}$ for that city. We refer to this learned policy as the Learned Constructor (LC). The policy can be used for planning in one of two modes: stochastic mode, in which each action $a_{t}$ is randomly sampled from the distribution $\pi_{\theta}(\cdot|s_{t})$, and greedy mode, in which each action is deterministically (“greedily”) chosen as the highest-scored under the policy, $a_{t}=\max_{a\in\mathcal{A}}\pi(a|s_{t})$.
The central component of the policy net is a graph attention net (GAT) Brody et al. (2021) which treats the city as a fully-connected graph on the nodes $\mathcal{N}$. Each node has an associated feature vector $\mathbf{x}_{i}$, and each edge a feature vector $\mathbf{e}_{ij}$, that contain information about location, demand, existing transit connections, and the street edge (if one exists) between $i$ and $j$. We note that a graph attention net operating on a fully-connected graph has close parallels to a Transformer model Vaswani et al. (2017), but unlike Transformers this architecture enables the use of edge features that describe known relationships between elements.
The GAT outputs node embeddings $\mathbf{y}_{i}$, which are operated on by one of two policy “heads”, depending on the timestep: $\textup{NN}_{ext}$ for choosing among extensions when the timestep $t$ is odd, and $\textup{NN}_{halt}$ for deciding whether to halt when $t$ is even. Figure 2 is a schematic illustration of the function of each of these three components in the route-planning MDP. The full details of the neural architecture are presented in Appendix A. We have released the code for our method and experiments to the public [^2] Available at https://www.cim.mcgill.ca/~mrl/hgrepo/transit_learning/.
@@ 4.1.1 Training
Following the work of Kool et al. (2019), we train the policy net using REINFORCE with baseline, a policy gradient method proposed by Williams (1992), and set $\gamma=1$. Since, by construction, the reward is 0 except at the final planning step, this implies the return $G_{t}$ is the same for all steps:
$G_{t}=\sum^{t_{end}}_{t^{\prime}}R_{t^{\prime}}=-C(\mathcal{C},\mathcal{R}_{final})$ (8)
The learning signal for each action $a_{t}$ is $G_{t}-b(s_{t})$, where the baseline $b(s_{t})$ is a learned quantity that estimates the average $G_{t}$ achieved by $\pi_{\theta}$ starting in state $s_{t}$. Differing from Kool et al. (2019), we compute the baseline at each step using a small Multi-Layer Perceptron (MLP) neural net separate from the policy network $\pi_{\theta}$. This baseline net is trained to predict $G_{t}$ based on cost weight parameter $\alpha$ and statistics of the city $\mathcal{C}$.
We train the policy on a dataset of synthetic cities. We sample a different $\alpha\sim[0,1]$ for each city, while holding $S,n,MIN,MAX$, and the constraint weight $\beta$ constant across training. The values of these parameters used are presented in Table 7. For each batch, a full MDP episode is run on the cities in the batch, $C(\mathcal{C},\mathcal{R})$ is computed across the batch, and back-propagation and weight updates are applied to both the policy net and the baseline net.
@@
  • 4-grid: Place $n$ nodes in a rectangular grid as close to square as possible. Add edges from each node to its horizontal and vertical neighbours.
  • 8-grid: The same as the above, but also add edges between diagonal neighbours.
  • Voronoi: Sample $m$ random 2D points, and compute their Voronoi diagram Fortune (1995). Take the shared vertices and edges of the resulting Voronoi cells as $\mathcal{N}$ and $\mathcal{E}_{s}$. $m$ is chosen so $|\mathcal{N}|=n$.
@@
To generate training data, we follow the general procedure used to generate the synthetic cities in the Mumford benchmark dataset Mumford (2013b). $\mathcal{N}$ is generated by sampling $n$ 2D points in a $30$ $\mathrm{km}$ by $30$ $\mathrm{km}$ square. The relationship between $n$ and $n_{e}=|\mathcal{E}_{s}|$ in Mumford’s cities approximately $n_{e}=3.6*n-34$, so we make this the desired number of edges. To generate $\mathcal{E}_{s}$, we start with the minimum spanning tree of the nodes in $\mathcal{N}$, and then add edges $(i,j,t_{ij})$ to this graph in ascending order of $t_{ij}=||(x_{i},y_{i})-(x_{j},y_{j})||_{2}$ until $n_{e}$ is reached. Finally, we generate the OD matrix $D$ by setting diagonal demands $D_{ii}=0$ and uniformly sampling off-diagonal elements $D_{ij}\sim[60,800]$.
We train our policies on a dataset of $7682^{15}=32,$ 768 synthetic cities. We made a 90:10 split of this dataset into training and validation sets. Training progresses in “epochs”, where one epoch corresponds to training the model on one episode over each city in the training set. After each epoch, the policy is run on the validation set, and its average cost is recorded. At the end of training, the parameters $\theta$ from the epoch with the lowest average validation cost are returned, giving the final policy $\pi_{\theta}$.
All neural net inputs are normalized so as to have unit variance and zero mean across the entire dataset during training. The scaling and shifting normalization parameters are saved as part of the model and applied to new data presented at test time.
@@ 4.2 Evolutionary Algorithm
Having trained a learned constructor policy, we may now use it as a low-level heuristic in metaheuristic algorithms. To demonstrate this and see whether it offers any benefit, we here describe an evolutionary algorithm with two mutation operators, and the modified algorithm which replaces one of these mutation operators with one that uses our learned constructor policy.
@@
The algorithm operates on a population of $B$ solutions $\mathcal{R}_{b}$, and functions by alternating stages of mutation and selection. In the mutation stage, the algorithm applies two mutation operators, “type 1” and “type 2”, each to a subset of the population chosen at random; if the resulting mutation has lower cost, it replaces its “parent”. This proceeds for $E$ steps. In the selection stage, solutions randomly either “die” or “reproduce”, with probabilities inversely related to their cost $C(\mathcal{C},\mathcal{R}_{b})$.
Each mutator begins by selecting, uniformly at random, a route $r$ in the solution and a terminal $i$ on that route. The type-1 mutator then selects a random node $j\neq i$ in $\mathcal{N}$, and replaces $r$ with the shortest path between $i$ and $j$. The probability of choosing each node $j$ is proportional to the amount of demand directly satisfied by $SP_{ij}$. The type-2 mutator chooses with probability $p_{d}$ to delete $i$ from $r$; otherwise, it adds a random node $j$ in $i$’s neighbourhood to $r$ (before $i$ if $i$ is the first node in $r$, and after $i$ if $i$ is the last node in $r$), making $j$ the new terminal. Following Nikolić and Teodorović (2013), we set the deletion probability $p_{d}=0.2$ in our experiments.
The initial population of solutions is constructed by copying a single initial solution, $\mathcal{R}_{0}$, $B$ times. In our prior work (Holliday and Dudek, 2023), we found that using the learned constructor in to plan $\mathcal{R}_{0}$ outperformed other methods, so we use this technique here. Specifically, we run the learned constructor in stochastic mode 100 times on a city, generating 100 distinct transit networks, and then choose the lowest-cost of these as $\mathcal{R}_{0}$. We refer to this procedure as “LC-100”.
The parameters of this algorithm are the population size $B$, the number of iterations of mutation-selection $IT$, and the number of mutations per mutation stage $E$. In addition to these, it takes as input the city $\mathcal{C}$ being planned over, the set of shortest paths through the city’s street graph SP, and the cost function $C$. The full procedure is given in 1. We refer to this algorithm as “EA”.
@@
2:Construct initial solution $\mathcal{R}_{0}$
3:$\mathcal{R}_{b}\leftarrow\mathcal{R}_{0}\;\forall\;b\in[1,B]$
@@
6:     // Explore around each solution
@@
10:                  $\mathcal{R}^{\prime}_{b}\leftarrow$ type_1_mod($\mathcal{R}_{b}$)
@@
12:                  $\mathcal{R}^{\prime}_{b}\leftarrow$ type_2_mod($\mathcal{R}_{b}$)
@@
15:         Randomly reorder solutions $\mathcal{R}_{b}$ in the population
16:     // Decide which solutions survive and reproduce
@@
22:         $O_{b}\leftarrow\frac{C_{max}-C_{b}}{C_{max}-C_{min}}$
23:         $s_{b}\sim\textup{Bernoulli}(1-e^{-O_{b}})$
24:         $p_{b}\leftarrow\frac{O_{b}s_{b}}{\sum_{b^{\prime}}O_{b^{\prime}}s_{b^{\prime}}}$
25:     if $\exists\;b\in[1,B]$ s.t. $s_{b}=1$ then
26:         for $b=1$ to $B$ do
27:              if $s_{b}=0$ then
28:                  $k\sim P(K)$, where $P(K=b^{\prime})=p_{b^{\prime}}$
29:                  $\mathcal{R}_{b}=\mathcal{R}_{k}$
30:return $\mathcal{R}_{\textup{best}}$
We note that Nikolić and Teodorović (2013) describe theirs as a “bee colony optimization” algorithm. As noted in Sörensen (2015), “bee colony optimization” is in fact a relabelling of the evolutionary-algorithm metaheuristic: the two are functionally identical. To avoid the profusion of unnecessary terminology, we here describe the algorithm as an evolutionary algorithm.
Our neural evolutionary algorithm NEA differs from EA only in that the type-1 mutator is replaced with a “neural mutator”. This mutator selects a route $r\in\mathcal{R}_{b}$ at random, deletes it from $\mathcal{R}_{b}$, and then rolls out our learned policy $\pi_{\theta}$ starting from $\mathcal{R}_{b}\setminus r$, generating one new route $r^{\prime}$ which is then added to $\mathcal{R}_{b}$. The algorithm is otherwise unchanged. We replace the type-1 mutator because its action space (replacing one route by a shortest path) is a subset of the action space of the neural mutator (replacing one route by a new route composed of shortest paths), while the type-2 mutator’s action space is quite different. In this way, $\pi_{\theta}$, which was trained as a construction policy, is applied to aid in an improvement method.

### v2

@@ 3.1 Markov Decision Process Formulation
A Markov Decision Process (MDP) is a formalism commonly used to define problems in RL (Sutton and Barto, 2018, Chapter 3). In an MDP, an agent interacts with an environment over a series of timesteps $t$, starting at $t=0$. At each timestep $t$, the environment is in a state $s_{t}\in\mathcal{S}$, and the agent observes the state and takes some action $a_{t}\in\mathcal{A}_{t}$, where $\mathcal{A}_{t}$ is the set of available actions at that timestep. The environment then transitions to a new state $s_{t+1}\in\mathcal{S}$ according to the state transition distribution $P(s_{t+1}|s_{t},a_{t})$, and the agent receives a numerical reward $R_{t}\in\mathbb{R}$ according to the reward distribution $P(R_{t}|s_{t},a_{t},s_{t+1})$. The agent chooses actions according to its policy $\pi(a_{t}|s_{t})$, which is a probability distribution over $\mathcal{A}_{t}$ given $s_{t}$. In RL, the goal is usually to learn a policy $\pi$ that maximizes the return $G_{t}$, defined as a time-discounted sum of rewards:
$G_{t}=\sum^{t_{\text{end}}}_{t^{\prime}=t}\gamma^{t^{\prime}-t}R_{t^{\prime}}$ (2)
Where $\gamma\in[0,1]$ is a parameter that discounts rewards farther in the future, and $t_{\text{end}}$ is the final time-step of the MDP. The sequence of states visited, actions taken, and rewards received from $t=0$ to $t_{\text{end}}$ constitutes one episode of the MDP.
We here describe the MDP we use to represent a construction approach to the transit network design problem. As shown in Equation 3, the state $s_{t}$ is composed of the set of finished routes $\mathcal{R}_{t}$, and an in-progress route $r_{t}$ which is currently being planned.
@@
The starting state is $s_{0}=(\mathcal{R}_{0}=\{\},r_{0}=[])$. At a high level, the MDP alternates at every timestep between two modes: on odd-numbered timesteps, the agent selects an extension to the route $r_{t}$ that it is currently planning; on even-numbered timesteps, the agent chooses whether or not to stop extending $r_{t}$ and add it to the set of finished routes.
On odd-numbered timesteps, the available actions are drawn from SP, the set of shortest paths between all pairs of nodes in $\mathcal{C}$. If $r_{t}=[]$, then:
$\mathcal{A}_{t}=\{a\;|\;a\in\textup{SP},|a|\leq MAX\}$ (4)
Otherwise, $\mathcal{A}_{t}=\text{EX}_{r_{t}}$, where $\text{EX}_{r_{t}}$ is the set of paths $a\in\textup{SP}$ that satisfy all of the following conditions:
@@
Once a path $a_{t}\in\mathcal{A}_{t}$ is chosen, $r_{t+1}$ is formed by appending $a_{t}$ to the beginning or end of $r_{t}$ as appropriate.
@@
$\mathcal{A}_{t}=\begin{cases}\{\textup{continue}\}&\text{if}|r_{t}|<MIN\\ \{\textup{halt}\}&\text{if}|r_{t}|=MAX\text{ or }|\text{EX}_{r_{t}}|=0\\ \{\textup{continue},\textup{halt}\}&\text{otherwise}\end{cases}$ (5)
If $a_{t}=\textup{halt}$, $r_{t}$ is added to $\mathcal{R}_{t}$ to get $\mathcal{R}_{t+1}=\mathcal{R}_{t}\cup\{r_{t}\}$, and $r_{t+1}=[]$ is a new empty route. If $a_{t}=\textup{continue}$, then $\mathcal{R}_{t+1}=\mathcal{R}_{t}$ and $r_{t+1}=r_{t}$.
The episode ends when the $S$-th route is added to $\mathcal{R}$: that is, if $|\mathcal{R}_{t+1}|=S$, the episode ends at timestep $t$. The output transit network is set to $\mathcal{R}=\mathcal{R}_{t+1}$, and the final reward is $R_{t_{\text{end}}}=-C(\mathcal{C},\mathcal{R})$. At all prior steps, $R_{t}=0$.
This MDP formalization imposes some helpful biases on the space of solutions. First, it requires any route connecting $i$ and $j$ to stop at all nodes along some path between $i$ and $j$, biasing planned routes towards covering more nodes. Second, it biases routes towards directness by forcing them to be composed of shortest paths. While an agent may construct arbitrarily indirect routes by choosing paths with length 2 at every step, this is unlikely because in a realistic street graph, the majority of paths in SP are longer than two nodes. Third, the alternation between deciding to continue or halt a route and deciding how to extend the route means that the probability of halting does not depend on how many different extensions are available; so a policy learned in environments with fewer extensions should generalize more easily to environments with more, and vice versa.
@@ 3.2 Cost Function
Where $\tau_{\mathcal{R}ij}$ is the time of the shortest transit trip from $i$ to $j$ given $\mathcal{R}$, including a time penalty $p_{T}$ for each transfer the rider must make between routes.
@@
To enforce the constraints on $\mathcal{R}$, we use a third term $C_{c}$, which is the fraction of node pairs $(i,j)$ with $D_{ij}>0$ that are not connected by $\mathcal{R}$, plus a measure of by how many stops $|r|>MAX$ or $|r|<MIN$ across all routes. The cost function is then:
@@
The weight $\alpha\in[0,1]$ controls the trade-off between passenger and operating costs, while $\beta$ is the penalty assigned for each constraint violation. $w_{p}$ and $w_{o}$ are re-scaling constants chosen so that $w_{p}C_{p}$ and $w_{o}C_{o}$ both vary roughly over the range $[0,1)$ for different $\mathcal{C}$ and $\mathcal{R}$; this is done so that $\alpha$ will properly balance the two, and to stabilize training of the GNN policy. The values used are $w_{p}=(\max_{i,j}T_{ij})^{-1}$ and $w_{o}=(S\max_{i,j}T_{ij})^{-1}$, where $T$ is an $n\times n$ matrix of shortest-path driving times between every node pair.
@@ 4.1 Learning to Construct a Network
In order to learn heuristics for transit network design, we first train a GNN policy $\pi_{\theta}(a|s)$, parameterized by $\theta$, to maximize the cumulative return $G_{t}$ on the construction MDP described in subsection 3.1. By then following this learned policy on the MDP for some city $\mathcal{C}$, stochastically sampling $a_{t}$ at each step with probabilities given by $\pi(\cdot|s_{t})$, we can obtain a transit network $\mathcal{R}$ for that city. We refer to this construction algorithm using the learned policy as “learned construction”. Because learned construction stochastically samples actions, we can run it multiple times to generate multiple networks for a city, and then pick the lowest-cost one. We will often do this by running learned construction 100 times and choosing the best network; we denote this procedure “LC-100”.
The central component of the policy net is a graph attention net (GAT) (Brody et al., 2021) which treats the city as a fully-connected graph on the nodes $\mathcal{N}$. Each node has an associated feature vector $\mathbf{x}_{i}$, and each edge a feature vector $\mathbf{e}_{ij}$, containing information about location, demand, existing transit connections, and the street edge (if one exists) between $i$ and $j$. We note that a graph attention net operating on a fully-connected graph has close parallels to a Transformer model (Vaswani et al., 2017), but unlike a Transformer, this architecture enables the use of edge features that describe known relationships between elements.
The GAT outputs node embeddings $\mathbf{y}_{i}$, which are operated on by one of two “head” neural nets, depending on the timestep: $\textup{NN}_{ext}$ for choosing among extensions when the timestep $t$ is odd, and $\textup{NN}_{halt}$ for deciding whether to halt when $t$ is even. Figure 2 is a schematic illustration of the function of each of these three components in the route-planning MDP. The full details of the neural architecture are presented in Appendix A. We have released the code for our method and experiments to the public [^2] Available at https://www.cim.mcgill.ca/~mrl/hgrepo/transit_learning/.
@@ 4.1.1 Training
Following the work of Kool et al. (2019), we train the policy net using REINFORCE with baseline, a policy gradient method proposed by Williams (1992), and set $\gamma=1$. Since, by construction, the reward $R_{t}=0$ except at $t_{\text{end}}$, this implies the return $G_{t}$ is the same for all steps:
$G_{t}=\sum^{t_{\text{end}}}_{t^{\prime}}R_{t^{\prime}}=-C(\mathcal{C},\mathcal{R})$ (9)
The learning signal for each action $a_{t}$ is $L_{t}=G_{t}-b(s_{t})$, where the baseline $b(s_{t})$ is a learned quantity that estimates the average $G_{t}$ achieved by $\pi_{\theta}$ starting in state $s_{t}$. Differing from Kool et al. (2019), we compute the baseline at each step using a small Multi-Layer Perceptron (MLP) neural net that is separate from the $\pi_{\theta}$. This baseline net is trained to predict $G_{t}$ based on the cost weight parameter $\alpha$ and statistics of the city $\mathcal{C}$.
We train the policy net on a dataset of synthetic cities. We sample a different $\alpha\sim[0,1]$ for each city, while holding $S,MIN,MAX$, and the constraint weight $\beta$ constant across training. The values of these parameters used are presented in Table 7. At each iteration, a full MDP episode is run on a “batch” of cities from the dataset, $C(\mathcal{C},\mathcal{R})$, $b(s_{t})$ and $L_{t}$ are computed across the batch, and back-propagation and parameter updates are applied to both the policy net (taking the gradient with respect to $-L_{t}\ln\pi_{\theta}$) and the baseline net (taking the gradient with respect to the mean squared error between $G_{t}$ and $b(s_{t})$).
@@
  • 4-grid: Place $n$ nodes in a rectangular grid as close to square as possible. Add edges between all horizontal and vertical neighbour nodes.
  • 8-grid: The same as 4-grid, but also add edges between diagonal neighbour nodes.
  • Voronoi: Sample $m$ random 2D points in a square, and compute their Voronoi diagram (Fortune, 1995). Take the shared vertices and edges of the resulting Voronoi cells as $\mathcal{N}$ and $\mathcal{E}_{s}$. $m$ is chosen so $|\mathcal{N}|=n$.
For each process except Voronoi, each edge in $\mathcal{E}_{s}$ is then deleted with probability $\rho$. If the resulting street graph is not strongly connected, it is discarded and the process is repeated.
@@
During training, we make a 90:10 split of this dataset into training and validation sets. Training progresses in “epochs”, where one epoch corresponds to running learned construction with the policy net on one MDP episode over each city in the training set. After each epoch, we run learned construction with the policy net on each city in the validation set, and the average cost of the resulting networks is recorded. At the end of training, the parameters $\theta$ from the epoch with the lowest average validation cost are returned, giving the final policy $\pi_{\theta}$.
All neural net inputs are normalized so as to have unit variance and zero mean across the entire dataset during training. The scaling and shifting normalization parameters are saved as part of the policy net and are applied to new data presented to $\pi_{\theta}$ after training.
@@ 4.2 Evolutionary Algorithm
We wish to consider whether our learned policy $\pi_{\theta}$ can be effective as a low-level heuristic in a metaheuristic algorithm. We test this by comparing a simple evolutionary algorithm to a variant of the algorithm where we replace one of its low-level heuristics with one that uses our learned policy. In this section we describe the baseline evolutionary algorithm and our variant of it.
@@
The algorithm operates on a population of $B$ solutions =$\mathcal{B}=\{\mathcal{R}_{b}|1\leq b\leq B\}$, and performs alternating stages of mutation and selection. In the mutation stage, the algorithm applies two “mutators” (ie. low-level heuristics), type 1 and type 2, to equally-sized subsets of the population chosen at random; if the mutated network $\mathcal{R}^{\prime}_{b}$ has lower cost than its “parent” $\mathcal{R}_{b}$, it replaces its parent in $\mathcal{B}$: $\mathcal{R}_{b}\leftarrow\mathcal{R}^{\prime}_{b}$. This is repeated $E$ times in the stage. Then, in the selection stage, solutions either “die” or “reproduce”, with probabilities inversely related to their cost $C(\mathcal{C},\mathcal{R}_{b})$. After $IT$ repetitions of mutation and selection, the algorithm returns the best network found over all iterations.
Both mutators begins by selecting, uniformly at random, a route $r$ in $\mathcal{R}_{b}$ and a terminal node $i$ on that route. The type-1 mutator then selects a random node $j\neq i$ in $\mathcal{N}$, and replaces $r$ with the shortest path between $i$ and $j$, $\text{SP}_{ij}$. The probability of choosing each node $j$ is proportional to the amount of demand directly satisfied by $\text{SP}_{ij}$. The type-2 mutator chooses with probability $p_{d}$ to delete $i$ from $r$; otherwise, it adds a random node $j$ in $i$’s street-graph neighbourhood to $r$ (before $i$ if $i$ is the first node in $r$, and after $i$ if $i$ is the last node in $r$), making $j$ the new terminal. Following Nikolić and Teodorović (2013), we set the deletion probability $p_{d}=0.2$ in our experiments.
The initial population of solutions is constructed by making $B$ copies of a single initial solution, $\mathcal{R}_{0}$. In our prior work (Holliday and Dudek, 2023), we found that using learned construction to plan $\mathcal{R}_{0}$ outperformed other methods, so we use this technique here. Specifically, we use LC-100 to generate $\mathcal{R}_{0}$.
The parameters of this algorithm are the population size $B$, the number of mutations per mutation stage $E$, and the number of iterations of mutation and selection stages $IT$. In addition to these, the algorithm takes as input the city $\mathcal{C}$ being planned over, the set of shortest paths through the city’s street graph SP, and the cost function $C$. The full procedure is given in Algorithm 1. We refer to this algorithm as “EA”.
@@
2:Construct initial network $\mathcal{R}_{0}\leftarrow\text{LC-100}(\mathcal{C},C)$
3:$\mathcal{R}_{b}\leftarrow\mathcal{R}_{0}\;\forall\;b\in\text{integers }1\text{through }B$
@@
6:     // Mutation stage
@@
10:                  $\mathcal{R}^{\prime}_{b}\leftarrow$ type_1_mut($\mathcal{R}_{b}$)
@@
12:                  $\mathcal{R}^{\prime}_{b}\leftarrow$ type_2_mut($\mathcal{R}_{b}$)
@@
15:         Randomly shuffle network indices $b$
16:     // Selection stage
@@
22:         // Select surviving networks
23:         $O_{b}\leftarrow\frac{C_{max}-C_{b}}{C_{max}-C_{min}}$
24:         $s_{b}\sim\textup{Bernoulli}(1-e^{-O_{b}})$
25:         // Set survivor reproduction probabilities
26:         $p_{b}\leftarrow\frac{O_{b}s_{b}}{\sum_{b^{\prime}}O_{b^{\prime}}s_{b^{\prime}}}$
27:     if $\exists\;b\in[1,B]$ s.t. $s_{b}=1$ then
28:         // Replace non-survivors with survivors’ offspring
29:         for $b=1$ to $B$ do
30:              if $s_{b}=0$ then
31:                  $k\sim P(K)$, where $P(K=b^{\prime})=p_{b^{\prime}}$
32:                  $\mathcal{R}_{b}=\mathcal{R}_{k}$
33:return $\mathcal{R}_{\textup{best}}$
We note that Nikolić and Teodorović (2013) describe theirs as a “bee colony optimization” algorithm. As noted in Sörensen (2015), “bee colony optimization” is merely a relabelling of the components of one kind of evolutionary algorithm. It is mathematically identical to this older, well-established metaheuristic. To avoid confusion and the spread of unnecessary terminology, we here describe the algorithm as an evolutionary algorithm.
Our neural evolutionary algorithm (NEA) differs from EA only in that the type-1 mutator is replaced with a “neural mutator”. This mutator selects a route $r\in\mathcal{R}_{b}$ at random, deletes it from $\mathcal{R}_{b}$, and then rolls out learned construction with $\pi_{\theta}$ starting from $\mathcal{R}_{b}\setminus r$. Because it starts from $\mathcal{R}_{b}\setminus r$, and $|\mathcal{R}_{b}\setminus r|=S$, this produces just one new route $r^{\prime}$, which then replaces $r$ in $\mathcal{R}_{b}$. The algorithm is otherwise unchanged. We replace the type-1 mutator because its space of changes (replacing one route by a shortest path) is similar to that of the neural mutator (replacing one route by a new route composed of shortest paths), while the type-2 mutator’s space of changes is quite different (lengthening or shortening a route by one node). In this way we leverage $\pi_{\theta}$, which was trained as a construction policy, to aid in an improvement method.